AnimeGhoul8863:

If the outliers are not included in the data set below, what is the mean of the data set? 42, 43, 46, 48, 57, 60, 96, 59, 38, 68, 29 47 48 49 52

AnimeGhoul8863:

@563blackghost

ThisGirlPretty:

I know this :D

AnimeGhoul8863:

really

ThisGirlPretty:

Yes

AnimeGhoul8863:

SWEET PLEASE HELP

ThisGirlPretty:

Step 1! First, you must add all the numbers up!

AnimeGhoul8863:

ok

AnimeGhoul8863:

so like 42+43+46ect

ThisGirlPretty:

42 + 43 + 46 + 48 + 57 + 60 + 96 + 59 + 38 + 68 + 29 =

AnimeGhoul8863:

ok give me a sec

ThisGirlPretty:

Take your time :D

AnimeGhoul8863:

586

AnimeGhoul8863:

ok whats next

ThisGirlPretty:

Good job!

ThisGirlPretty:

Next, we have divide the amount of numbers that you multiplied for an example like 42 + 43 + 46 + 48 + 57 + 60 + 96 + 59 + 38 + 68 + 29

ThisGirlPretty:

How many numbers are there?

AnimeGhoul8863:

so 586/11=53.272727272727

ThisGirlPretty:

Yeah you're right

Moon:

it says to remove outliers so I would recommend removing 96 before calculating the mean

ThisGirlPretty:

She's right I miss read that.

AnimeGhoul8863:

ok

AnimeGhoul8863:

58.6

ronndjaia:

d

AnimeGhoul8863:

idid do d and got it wrong

ThisGirlPretty:

Hmm... I got 44.5454545454545....@Vocaloid whenever you get the chance >.<

ThisGirlPretty:

Let me try one more thing @AnimeGhoul8863

AnimeGhoul8863:

i got 49

AnimeGhoul8863:

so imthinking c

ThisGirlPretty:

That should be it >.<. Can you show me how you got it please?

ThisGirlPretty:

I wanna learn also >.<

AnimeGhoul8863:

42 + 43 + 46 + 48 + 57 + 60 + 59 + 38 + 68 + 29= 490 divide by 10 =49

ronndjaia:

the answer has to be c 49

AnimeGhoul8863:

thx

ThisGirlPretty:

Now I know what I messed up on, thank you haha :D!

AnimeGhoul8863:

thank u too lol

ThisGirlPretty:

No problem :)